# M2SA: Multimodal and Multilingual Model for Sentiment Analysis of Tweets

### v2

@@ 3.3.   Dataset
The dataset contains the following languages: Arabic-ar (Nakov et al., 2013b; Yin et al., 2021), Bulgarian-bg (Mozetič et al., 2016), Bosnian-bs (Mozetič et al., 2016), Danish-da (Pauli et al., 2021), German-de (Rei et al., 2016), English-en (Nakov et al., 2013c; Ghosh et al., 2015; Nakov et al., 2013d, a; Vanzo et al., 2014; Castellucci et al., 2015; Tayebi Arasteh et al., 2021), Spanish-es (Adrián, 2016-05-25; Santamaría et al., 2022; Agüero-Torales et al., 2021; Villena-Román and Garcıa-Morera, 2013; Román et al., 2015; Vilares et al., 2015; Montejo-Ráez and Díaz-Galiano, 2016; Cámara et al., 2018; Díaz-Galiano et al., 2019; García-Vegaa et al., 2020), French-fr (Vukotić et al., 2015), Croatian-hr (Babić et al., 2021), Hungarian-hu (Mozetič et al., 2016), Italian-it (Moctezuma et al., 2016) , Maltese-mt (Cortis and Davis, 2021), Polish-pl (Mozetič et al., 2016), Portuguese-pt (Patrick et al., 2022), Russian-ru (Mozetič et al., 2016), Serbian-sr (Ljajić and Marovac, 2018), Swedish-sv (Mozetič et al., 2016), Turkish-tr (Mutlu and Özgür, 2022), Chinese-zh (Yin et al., 2021), Latvian-lv (Muischnek and Müürisep, 2018) and Albanian-sq (Mozetič et al., 2016). Languages with more data points, such as German, Spanish, English, Italian, Arabic, and Polish, possess dataset instances exceeding 10,000, whereas other languages exhibit 5,000 or fewer instances of text and images. The dataset exhibits an average token count ranging from 4.25 to 5.94 words, separating each token by a space. One observable pattern is that tweets classified as positive tend to be more likely to be accompanied by images than other categories. The diagram also indicates an imbalance in the datasets across the languages.
@@ 8.   Bibliographical References
  • Muischnek and Müürisep (2018) K Muischnek and K Müürisep. 2018. Latvian tweet corpus and investigation of sentiment analysis for latvian. In Human Language Technologies–The Baltic Perspective: Proceedings of the Eighth International Conference Baltic HLT 2018, volume 307, page 112. IOS Press.

### v3

@@ 3.3.   Dataset
The dataset contains the following languages: Arabic-ar (Nakov et al., 2013b; Yin et al., 2021), Bulgarian-bg (Mozetič et al., 2016), Bosnian-bs (Mozetič et al., 2016), Danish-da (Pauli et al., 2021), German-de (Rei et al., 2016), English-en (Nakov et al., 2013c; Ghosh et al., 2015; Nakov et al., 2013d, a; Vanzo et al., 2014; Castellucci et al., 2015; Tayebi Arasteh et al., 2021), Spanish-es (Adrián, 2016-05-25; Santamaría et al., 2022; Agüero-Torales et al., 2021; Villena-Román and Garcıa-Morera, 2013; Román et al., 2015; Vilares et al., 2015; Montejo-Ráez and Díaz-Galiano, 2016; Cámara et al., 2018; Díaz-Galiano et al., 2019; García-Vegaa et al., 2020), French-fr (Vukotić et al., 2015), Croatian-hr (Babić et al., 2021), Hungarian-hu (Mozetič et al., 2016), Italian-it (Moctezuma et al., 2016) , Maltese-mt (Cortis and Davis, 2021), Polish-pl (Mozetič et al., 2016), Portuguese-pt (Patrick et al., 2022), Russian-ru (Mozetič et al., 2016), Serbian-sr (Ljajić and Marovac, 2018), Swedish-sv (Mozetič et al., 2016), Turkish-tr (Mutlu and Özgür, 2022), Chinese-zh (Yin et al., 2021), Latvian-lv (Pinnis, 2018) and Albanian-sq (Mozetič et al., 2016). Languages with more data points, such as German, Spanish, English, Italian, Arabic, and Polish, possess dataset instances exceeding 10,000, whereas other languages exhibit 5,000 or fewer instances of text and images. The dataset exhibits an average token count ranging from 4.25 to 5.94 words, separating each token by a space. One observable pattern is that tweets classified as positive tend to be more likely to be accompanied by images than other categories. The diagram also indicates an imbalance in the datasets across the languages.
@@ 8.   Bibliographical References
  • Pinnis (2018) Marcis Pinnis. 2018. Latvian tweet corpus and investigation of sentiment analysis for latvian. In Human Language Technologies - The Baltic Perspective - Proceedings of the Eighth International Conference Baltic HLT 2018, Tartu, Estonia, 27-29 September 2018, volume 307 of Frontiers in Artificial Intelligence and Applications, pages 112–119. IOS Press.